If a certain object has a density of 0.4 g/cm3, what is its specific gravity?


1
Expert's answer
2022-03-19T02:45:43-0400

p = 0.4 g/cm^3=400kg/m^3

g = 9.8 m/s^2

Y - ?

Solving:

Y= p*g

Y = 400kg/m^3 * 9.8m/s^2= 3920 H/m^3

Answer: 3920 H/m^3